OpenStudy (anonymous):

A carpenter cut four lengths of wood. The lengths were 36 3/4 in., 36 3/8 in., 37 1/2 in., and z in. If the mean of the lengths is 36 5/8 in., what was the length of the fourth piece of wood the carpenter cut? A: 36 5/8 in B: 35 7/8 in C: 36 7/8 in D: 37 1/4 in

OpenStudy (anonymous):

the average is the sum of the four lengths divided by 4

OpenStudy (anonymous):

Do you know the answer?

OpenStudy (anonymous):

yes, but the objective is to help you to figure it out

OpenStudy (nurali):

(36.75 + 36.375 + 37.5 + z)/4 = 36.625 36.75 + 36.375 + 37.5 + z = 146.5 z = 146.5 - 36.75 - 36.375 - 37.5 = 35.875 = 35 7/8 in

OpenStudy (anonymous):

Thanks a lot!

OpenStudy (nurali):

Anytime

OpenStudy (anonymous):

Just did. :p